How many quarts of a 10% antifreeze solution must be mixed with 16 quartz of a 40% antifreeze solution to make a 30% solution?
Answer by richwmiller(17219)

30%=.3
10%=.1
40%=.4
0.1*x=10% antifreeze solution
0.4*16=16 quartz of a 40% antifreeze solution
0.3(16+x)=30% solution
0.1*x+0.4*16=0.3(16+x)
0.1*x+6.4=4.8+0.3x
0.1x-0.3x=4.8-6.4
-0.2x=-1.6
x=8 quarts of 10% antifreeze solution

check our math
0.1*8+0.4*16=0.3(16+8)
0.8+6.4=0.3(24)
7.2=7.2
ok